OpenStudy (anonymous):

A cell phone company charges a monthly fee of $20 for the first 1000 text messages and 40 cents for each additional text message. Miriam’s bill for text messages for the month of June is $49.20. How many text messages did she send that month?

OpenStudy (kittiwitti1):

Equation:\[$20+40n, n\text{ being amount of additional text messages}\]

OpenStudy (kittiwitti1):

\[20+40n=49.20\rightarrow40n=29.20\rightarrow n=\frac{29.20}{40}=?\]\[\text{since the 20 is for the first 1000 text messages add 1000 to the result you got above.}\]\[n+1000\rightarrow\frac{29.20}{40}+1000\]